What is today’s equivalent present value of $200,000 at the interest rate of 8% per year compounded annually, 10 years from today?

Please show full steps
Answer #1

Answer

By coumpounding table formula:

PV=F*(P|F,i,n)

F=$200000

i=8%

n=10 years

PV=200000*(P|F,8%,10)

=200000*0.463193488

=92638.6976

The value is $92638.6976

======

Or

By formula

The amount is $92638.69762